A rectangle plot of land is 50 m wide. The cost fencing the plot at the rate of Rs 18 per meter is rs 4680 find.
1) The length of the plot.
2) The cost of leveing the plot at rate of Rs 7.6 square meter.

Answer by mananth(16946)

let length be x
width = 50 m
..
Cost of fensing = 4680
rate = 18
perimeter = 4680/18=260
Perimeter = 2*(x+50)= 260
divide by 2
x+50 =130
x= 130-50
x= 80 m the length.
..
Area = L*W
Area = 80*50
Area = 4000 m^2
Rate of leveling = Rs. 7.6/m
Total Cost = 4000*7.6= Rs. 30,400
m.ananth@hotmail.ca